OpenStudy (nightwielder):

Medal ofcourse. Jodi pours herself a cold soda that has an initial temperature of 38 F and immediately went outside to sunbathe where the temperature was a steady 95 F. After 7 minutes the temperature of the soda was 46 F. Jodi had to run back into the house to answer the phone. What is the expected temperature of the soda after an additional 10 minutes?

OpenStudy (anonymous):

easy way or hard way?

OpenStudy (anonymous):

ah hell lets to it the easy way because... it is easy

OpenStudy (nightwielder):

\[T(t)=S+(T _{0}-S)e ^{kt}\]

OpenStudy (perl):

you can use an exponential model, you are given two points

OpenStudy (nightwielder):

46=95+(-57)e^7k

OpenStudy (anonymous):

that is the ridiculous way lets make it simple start with the temp at 95 and the outside temp at 38 we work with the differnece in temps, which is \(95-38=57\)

OpenStudy (nightwielder):

ok

OpenStudy (nightwielder):

I was using newtons law of cooling

OpenStudy (anonymous):

in ten minutes the difference in temp is \(95-46=49\) yeah i know

OpenStudy (nightwielder):

no starts with 38

OpenStudy (nightwielder):

it starts cold

OpenStudy (anonymous):

ok lets back up

OpenStudy (anonymous):

outside temp is 95 beer is 38 difference is \(95-38=57\)

OpenStudy (perl):

i did y = 38 * e^(k*t) 46 = 38 * e^(k*t) , solve for k

OpenStudy (anonymous):

in ten minutes, outside temp is still 95, beer is 46, difference is \(95-46=57\)

OpenStudy (nightwielder):

sattelite, u r the only person i know that can turn soda into beer. :)

OpenStudy (nightwielder):

I want to see what ur trying to do.

OpenStudy (anonymous):

ratio is \[\frac{57}{46}\] in ten minutes in another ten minutes it will be \((\frac{57}{46})^2\)

OpenStudy (nightwielder):

57/46 in 7 mins

OpenStudy (anonymous):

take \[57\times \left(\frac{57}{46}\right)^2\] to get the new difference

OpenStudy (perl):

I got y = 38 * e^( .0191 * t )

OpenStudy (anonymous):

then subtract that from \(95\)

OpenStudy (anonymous):

@perl the 38 is wrong you work with the differences in the temperatures, not the actual temperature

OpenStudy (nightwielder):

i had said 10 minutes. i meant 7 minutes. its alright if u want to stop working on the problem. I can work it out by myself.

OpenStudy (anonymous):

ok if it is 7 minutes, then we have to do something different

OpenStudy (nightwielder):

After 7* minutes the temperature of the soda was 46 F.

OpenStudy (nightwielder):

that's why i did this to solve for k: 46=95+(-57)e^7k

OpenStudy (anonymous):

ok

OpenStudy (nightwielder):

so... -49=-57e^7k

OpenStudy (nightwielder):

49/57=e^7k ln(49/57)=7k

OpenStudy (nightwielder):

[ln(49/57)]/7=k

OpenStudy (perl):

ok i was using the wrong model.

OpenStudy (anonymous):

looks like you got it to me, once you have the model, replace \(t\) by \(17\)

OpenStudy (perl):

temperature change is proportional to the difference in temp. not the actual temperature?

OpenStudy (nightwielder):

k=-0.02160442424

OpenStudy (anonymous):

@perl yes, the difference decays to zero

OpenStudy (nightwielder):

T(17)=95+(-57)e^17k T(17)=95-39.4792830925

OpenStudy (nightwielder):

T(17)=55.5207169075

OpenStudy (perl):

you changed your question to 7 minutes

OpenStudy (perl):

you can edit your question

OpenStudy (nightwielder):

ok

OpenStudy (perl):

Jodi pours herself a cold soda that has an initial temperature of 38 F and immediately went outside to sunbathe where the temperature was a steady 95 F. After *7* minutes the temperature of the soda was 46 F. Jodi had to run back into the house to answer the phone. What is the expected temperature of the soda after an additional 10 minutes?

OpenStudy (nightwielder):

56 is not workingn -_-

OpenStudy (perl):

one sec, checking

OpenStudy (perl):

T(t) = Ta + ( To - Ta) e^(-kt)

OpenStudy (perl):

T(t) = 95 + (38 - 95) e^(-kt)

OpenStudy (nightwielder):

i dont have -kt in the equation i was given

OpenStudy (nightwielder):

but it seems that it wont matter

OpenStudy (perl):

right, it wont matter

OpenStudy (perl):

k is usually considered positive

OpenStudy (perl):

so in decay problems, we put a negative in front of k

OpenStudy (nightwielder):

ok

OpenStudy (perl):

did you get k = ln (49/57) / 7

OpenStudy (perl):

T(t) = 95 -57 e^(-.021604 * t )

OpenStudy (perl):

T(17) = 55.5

OpenStudy (nightwielder):

I had to round to the whole for the previous questions so I did the same here. I trid 56 snd 55. It turs out it wanted it rounded to the tenth. 55.5 is CORRECT!

OpenStudy (nightwielder):

Thanks so much @perl . Could u medal @satellite73 and ill medal u.

OpenStudy (perl):

sure

OpenStudy (nightwielder):

I could have just solved this myself but I came here. The power of laziness. :)

OpenStudy (perl):

:D

OpenStudy (nightwielder):

XD